# Efficient Approximate Degenerate Ordered Statistics Decoding for Quantum Codes via Reliable Subset Reduction

### v2

@@ III-B OSD4-$w$
Figure 1: A tree with nodes consisting of all bit strings of length 4 and with at most weight 2. The root node is $0$. A child node is obtained by flipping one of the zero bits from its parent node that come after the last bit that is one.
III-C Complexity of OSD4
We analyze the complexity of OSD based on 00-ADD in the following. The complexity of OSD4-0 is dominated by the Gaussian elimination step, which is then $O(2nm^{2})=O(n^{3})$, assuming $m=O(n)$. For OSD4-$w$ with $w>0$, the total number of possible flips is $\sum_{j=0}^{w}\binom{n+k}{j}$, which can be expressed as $O((n+k)^{w})$ or simply $O(n^{w})$.
@@ III-C Complexity of OSD4
Algorithm 3 HRSR
Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, hard-decision vector ${\mathbf{e}}\in\{0,1\}^{2n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, and soft reliability threshold $\theta$.
Output: shortened variables $\tilde{{\mathbf{s}}}$, $\tilde{{\mathbf{H}}}$, ${\mathbf{e}}^{H}$, and $\sigma$.
Steps:
1:Use $\bm{\eta}$, ${{\mathbf{q}}_{i}}_{i=1}^{n}$, and $\theta$ to identify the highly reliable bits in ${\mathbf{e}}$ based on Definition 7. Assume that there are $v$ such highly reliable bits.
2:Let $\sigma$ be a column permutation such that $\sigma({\mathbf{e}})=\begin{bmatrix}{\mathbf{e}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}$, where ${\mathbf{e}}^{H}$ is the highly reliable subset of $v$ bits, and ${\mathbf{e}}^{\prime}$ is the remaining $2n-v$ bits. Let $\sigma(\hat{{\mathbf{e}}})=\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}$ be a permuted error estimate, which contains the highly reliable bits in ${\mathbf{e}}^{H}$.
3:Let $\tau$ be a row permutation such that $\tau(\sigma({\mathbf{H}}\Lambda))=\begin{bmatrix}\tilde{{\mathbf{H}}}&{\mathbf{B}}\\ \mathbf{0}&{\mathbf{C}}\\ \end{bmatrix}$ for $\tilde{{\mathbf{H}}}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ with $m^{\prime}\leq m$. The $m-m^{\prime}$ rows $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$ correspond to parity checks related only to the highly reliable bits. Assume that $\tau({\mathbf{s}})=\begin{bmatrix}{\mathbf{s}}^{\prime\prime}\\ {\mathbf{s}}^{H}\end{bmatrix}$, where ${\mathbf{s}}^{\prime\prime}\in\{0,1\}^{m^{\prime}\times 1}$, and ${\mathbf{s}}^{H}\in\{0,1\}^{(m-m^{\prime})\times 1}$ are the syndrome bits corresponding to $\begin{bmatrix}\mathbf{0}&{\mathbf{C}}\end{bmatrix}$.
4:Verify the parity check condition related to the highly reliable bits.
if ${\mathbf{s}}^{H}\neq{\mathbf{C}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$ then return “Failure";
else we have a new parity-check condition
     where $\tilde{{\mathbf{s}}}={\mathbf{s}}^{\prime\prime}-{\mathbf{B}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$.
     if (15) is solvable then return $\tilde{{\mathbf{s}}}$, $\tilde{{\mathbf{H}}}$, ${\mathbf{e}}^{H}$, and $\sigma$;
     else return “Failure".
     end if
end if
IV Highly reliable subset reduction
As we increase the order $w$ of OSD, we can expect an improvement in the decoder’s performance. In particular, when $w=n+k$ and $\delta=n$, OSD4-$(n+k)$ using $n$-ADD becomes a maximum likelihood decoder. However, increasing the order $w$ also leads to an increase in time complexity.
To address this issue, we propose an approach to reduce the problem size for OSD. Recall that OSD is used to solve a linear system with $2n$ binary variables and $n-k$ constraints. If a subset of binary variables is known, the problem size–and consequently the computational complexity–can be reduced. The number of binary variables in the reduced linear system is referred to as the effective length for OSD, which corresponds to the length of the remaining binary error variable vector. We have the following definition of highly reliable coordinates.
Definition 7.
An error bit ${\mathbf{e}}_{i}^{a}$, where $i\in{1,\dots,n}$ and $a\in{X,Z}$, is considered highly reliable if its hard-decision reliability equals ${\mathrm{T}}$ or ${\mathrm{T}}+1$, and its soft reliability $\phi^{a}(i)$ satisfies $\phi^{a}(i)\geq\theta$, where $\theta$ is a predetermined reliability threshold.
An error variable has hard reliability ${\mathrm{T}}$ if BP has a constant output at all iterations; a hard reliability of ${\mathrm{T}}+1$ means the corresponding error variable is the identity during the BP process after being initialized to the identity. The soft reliability threshold $\theta$ is chosen close to one to ensure that a highly reliable coordinate not only exhibits an almost constant hard-decision history but also has a high soft reliability value.
These highly reliable variables are crucial in the OSD process. Identifying these variable nodes allows us to greatly reduce the computation burden of OSD.
These highly reliable parts provide outcomes that can be trusted and will remain unchanged throughout the OSD4 procedure with high probability. Consequently, we propose to exclude these highly reliable parts from the parity check relations. By doing so, we can avoid unnecessary calculations and prioritize our focus on the unreliable part. Our highly reliable subset reduction (HRSR) algorithm is given in Algorithm 3.
Equation (15) follows from the parity-check condition (2) with
$\begin{bmatrix}{\mathbf{s}}^{\prime\prime}\\ {\mathbf{s}}^{H}\end{bmatrix}=$ $\tau(\sigma({\mathbf{H}}\Lambda))\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}^{\top}=$ $\begin{bmatrix}\tilde{{\mathbf{H}}}\big{(}\hat{{\mathbf{e}}}^{\prime}\big{)}^{\top}+{\mathbf{B}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}\\ {\mathbf{C}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}\end{bmatrix}$
and the conditions ${\mathbf{s}}^{H}={\mathbf{C}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$ and $\tilde{{\mathbf{s}}}={\mathbf{s}}^{\prime\prime}-{\mathbf{B}}\big{(}{\mathbf{e}}^{H}\big{)}^{\top}$.
If HRSR returns a reduced linear system as in (15), we then use OSD4-$w$ to solve it with hard-decision BP output ${\mathbf{e}}^{\prime}$ and corresponding reliabilities $\bm{\eta}$, $\phi^{X}$, and $\phi^{Z}$. Suppose that OSD4-$w$ returns $\hat{{\mathbf{e}}}^{\prime}$. Then the error estimate is $\sigma^{-1}\left(\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}\right)$.
Reducing the problem size may cause the reduced check matrix to lose the ability to represent all valid codewords if its dimension becomes too small. To address this, in Step 4 of Algorithm 3, we verify that the removed highly reliable bits align with the corresponding syndrome bits. If they do not, the algorithm reports a failure.
This method effectively reduces the overhead in OSD, especially at low error rates as will be shown in the simulation section. Furthermore, as the effective length in OSD is reduced, we can apply higher order OSD if only a fixed amount of computation is allowed. Consider, for instance, the case where the number of error candidates to be tested in OSD is $\Gamma$. For the original decoding problem, OSD4-$2$ has
@@ IV Highly reliable subset reduction
Using the HRSR algorithm for $v$ highly reliable bits, we have a reduced system of linear equations with $2n-v$ binary variables. Assume the rank of the reduced system is $n-k-(m-m^{\prime})$. Then we can apply OSD4-$w$ under the constraint of $\Gamma$ candidates, with $w$ given by
$w=\operatorname*{arg\,max}_{x}\sum_{i=0}^{x}\binom{n+k-v+(m-m^{\prime})}{i}\leq\Gamma.$ (17)
Note that even though the number of tested candidates is fixed, the complexity of OSD following the HRSR algorithm is lower due to the reduced effective length.
@@ V Approximate degenerate OSD
In this section, we discuss how degeneracy in quantum codes helps reduce the computation of OSD. Recall that in OSD-$w$, $\sum_{i=0}^{w}\binom{n+k}{i}$ candidates are generated, and the one that is optimal with respect to $\delta$-ADD is chosen as the output. As shown in Lemma 6, a new candidate can be generated from an error candidate by flipping a bit in the reliable subset and adding a certain vector to the unreliable subset. This corresponds to adding a logical operator in the normalizer group to the error candidate.
By examining these logical operators, we have the following degenerate error candidate pruning conditions.
@@
Consider an $[[n,k,d]]$ stabilizer code with a check matrix given by $\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ after Gaussian elimination and a column permutation as in (7), where ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$ are associated with the reliable bits of OSD4-0 output.
@@
    When using $n$-ADD, only the reliable bits corresponding to nontrivial logical operators need to be considered in OSD4-$w$.
@@
    If each logical operator induced by flipping a reliable bit of OSD4-0 output is a stabilizer, then OSD4-$w$ is equivalent to OSD4-0 for any $w$.
@@
1) In $n$-ADD, different stabilizer cosets are compared. Therefore, multiplying a stabilizer by an error candidate does not create a new nontrivial candidate.
2) The statement is straightforward, as flipping these reliable bits results in only degenerate errors to the OSD4-0 output.
Exploiting Theorem 8-1) with $n$-ADD makes OSD4-$w$ more effective. By incorporating with the HRSR algorithm, we can work on a shortened check matrix $\tilde{{\mathbf{H}}}$ and higher order of OSD can be executed. Consequently, (17) can be improved to $w=\operatorname*{arg\,max}_{x}\sum_{i=0}^{x}\binom{u}{i}\leq\Gamma,$ where $u\leq n+k-v+(m-m^{\prime})$ is the number of columns in the shortened check matrix corresponding to nontrivial logical operators.
Proposition 9.
Checking whether a logical operator induced by flipping a reliable bit of the OSD4-0 output is a stabilizer can be done in $O((n-k)k)$ time.
Proof.
Consider an $[[n,k,d]]$ stabilizer code with a $m\times{2n}$ check matrix ${\mathbf{H}}$ and a $2k\times 2n$ binary logical matrix ${\mathbf{L}}$. During the process of OSD4-0, ${\mathbf{H}}$ leads to $\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ as in (7) using column permutations $\pi,\mu$ and row operations $R$, where ${\mathbf{A}}\in\{0,1\}^{(n-k)\times(n+k)}$ are associated with the reliable bits of OSD4-0 output.
By (13), flipping the $j$-th reliable bit corresponds to adding the vector ${\mathbf{g}}=\begin{bmatrix}{\mathbf{A}}_{j}^{\top}&0\cdots 0&1&0\cdots 0\end{bmatrix}\in\{0,1\}^{2n}$, where ${\mathbf{A}}_{j}$ is the $j$-th column of ${\mathbf{A}}$. The vector ${\mathbf{g}}$ corresponds to a stabilizer if $\mu(\pi({\mathbf{L}}\Lambda)){\mathbf{g}}^{\top}=\mathbf{0}$, and otherwise, it represents a nontrivial logical operator. Since ${\mathbf{A}}_{j}$ is of dimension $n-k$, this computation takes $O((n-k)k)$ time.
The computational complexity of verifying such a bit-flipping event is $O((n-k)k)$, which implies that checking all the columns of $A$ has a computational complexity of roughly the same order as Gaussian elimination. This approach is both effective and relatively efficient for implementing high-order OSD.
However, the complexity of $n$-ADD is not practically feasible. We may use $\delta$-ADD by choosing $\delta$ as the highest weight of a set of low-weight stabilizer generators. For the case of a rotated surface code [25], the stabilizer generators are of weight $2$ or $4$, and the next lowest stabilizer weight is $6$. Thus, $n$-ADD can be well approximated by $4$-ADD at error rates lower than $1\%$. Assume that OSD4444-0 generates a list of error candidates $\mathcal{E}$. Then, $4$-ADD outputs:
@@
For codes with algebraic structures, the first few lowest-weight terms can be derived, but determining the complete weight enumerator for a general code is an NP-hard problem. As an alternative, we propose selecting a set of $m\geq n-k$ low-weight stabilizer generators and calculating the weight distribution of (${m\choose c}$ stabilizers, where $c$ is a constant independent of $n$. This method provides a good approximation of the dominant terms for practical purposes. To maintain an overall complexity of $O(n^{3})$, the value of (${m\choose c}$ should be chosen accordingly.
To achieve optimal execution complexity, 00-ADD is often used. Theorem 8 suggests focusing on the columns of $A$ corresponding to nontrivial logical operators when $n$-ADD is used. However, in the scenario of 00-ADD, a stabilizer may influence the output of $\text{OSD}_{4}$-$w$ since multiplying a stabilizer can reduce the weight of the error, potentially providing a better candidate in 00-ADD, even though the error coset remains unchanged. Thus, $\text{OSD}_{4}$-$w$ cannot simply flip bits corresponding to those nontrivial logical operators as in the case of Theorem 8-1).
However, Theorem 8-2) still works and we may use Proposition 9 to check this condition.
In the following we provide a weaker sufficient but operationally more efficient condition, for degenerate error candidate pruning.
@@
Let $\tilde{S}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ be a shortened check matrix by the HRSR algorithm. Suppose that $\tilde{S}$ can be transformed into $\begin{bmatrix}{\mathbf{I}}&\tilde{{\mathbf{A}}}\end{bmatrix}$ after Gaussian elimination and a column permutation, where $\tilde{{\mathbf{A}}}$ is associated with the reliable bits. If all the columns of $\tilde{{\mathbf{A}}}$ are of weight less than $d-1$, then OSD4-$w$ on the corresponding reduced linear system is equivalent to OSD4-0 for any $w$.
The operational degenerate error pruning condition in Corollary 10 provides a more operationally efficient method to determine whether order-$w$ OSD for $w>0$ is nontrivial by evaluating the weight of each column. This can be done in $O((n-k)\times(n+k))$.
If there are many highly reliable bits, the number of stabilizers associated with these bits would be high. Consequently, the shortened check matrix $\tilde{\mathbf{S}}$ from the HRSR algorithm will have fewer rows and lower column weights. Therefore, Corollary 10 can be applied with high probability when the error rate is low.
We combine all the aforementioned techniques into the approximate degenerate OSD (ADOSD) algorithm, as detailed in Algorithm 4, using 00-ADD. Note that if decoding performance is of higher priority, $\delta$-ADD can be utilized, and Proposition 9 can be integrated into Algorithm 4.
@@
Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank $n-k$, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, BP output distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, soft reliability threshold $\theta$, hard-decision vector ${\mathbf{e}}\in\{0,1\}^{2n}$, maximum number of OSD flips $\Gamma$, and backup OSD order $w$.
Output: A valid error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$.
@@
Apply the HRSR algorithm.
if  HRSR returns failure, then
     use OSD4-$w$ to find an error estimate $\hat{{\mathbf{e}}}\in\{0,1\}^{2n}$ for the original problem, and return $\hat{{\mathbf{e}}}$;
else HRSR returns ${\mathbf{s}}^{\prime}$, $\tilde{{\mathbf{H}}}$, ${\mathbf{e}}^{H}$, and $\sigma$.
     if all the columns of $\tilde{{\mathbf{H}}}$ in reduced row echelon form have weight less than $d-1$then $w\leftarrow 0$;
@@
     Use OSD4-$w$ on ${\mathbf{s}}^{\prime}$, $\tilde{S}$, and the given data to find an error estimate $\hat{{\mathbf{e}}}^{\prime}\in\{0,1\}^{2n-v}$, and return $\sigma^{-1}\left(\begin{bmatrix}\hat{{\mathbf{e}}}^{\prime}&{\mathbf{e}}^{H}\end{bmatrix}\right)$.
@@
VI Simulation results
We simulate the performance of the MBP4+ADOSD4 scheme on LCS codes [60], GHP codes [43, 44], BB codes [56], and various 2D topological codes, including toric, rotated surface, color codes and twisted XZZX codes [23, 24, 25, 26, 27, 59, 29] (see [64, Table I]).
We use the MBP4 decoder introduced in [32] as a predecoder for OSD. MBP4 features a parameter $\alpha$ that controls the step size of message updates during BP iterations: when $\alpha<1$, the step size is enlarged; when $\alpha>1$, the step size is reduced. For $\alpha=1$, MBP4 is equivalent to the refined BP4 in [13]. We will specify the value of $\alpha$ only when $\alpha\neq 1$ is used. The optimization of $\alpha$ is discussed in Section VI-A.
The maximum number of BP iterations, ${\mathrm{T}}$, is set to $100$ unless otherwise specified. Furthermore, our MBP4+ADOSD4 scheme demonstrates robustness to variations in the maximum number of iterations, as discussed in Section VI-B. The soft reliability threshold is set to $\theta=0.999995$ and its selection is discussed in Section VI-C. In ADOSD4 simulations, $\Gamma$ is chosen according to (16) to enable comparison with OSD-2.
In Section VI-D, we compare the proposed BP+OSD schemes with various reliability measures using GHP codes. We then evaluate decoding performance on BB codes and topological codes in Sections VI-E, and VI-F. Finally, Section VI-G analyzes the execution time of these schemes, demonstrating the efficiency of ADOSD.

### v3

@@ IV-A OSD4-$w$
Figure 1: A tree with nodes consisting of all bit strings of length 4 and with at most weight 2. The root node is $00$. A child node is obtained by flipping one of the zero bits from its parent node that come after the last bit that is one.
IV-B Complexity of OSD4
We analyze the complexity of OSD based on 0-ADD in the following. The complexity of OSD4-0 is dominated by the Gaussian elimination step, which is then $O(2nm^{2})=O(n^{3})$, assuming $m=O(n)$. For OSD4-$w$ with $w>0$, the total number of possible flips is $\sum_{j=0}^{w}\binom{n+k}{j}$, which can be expressed as $O((n+k)^{w})$ or simply $O(n^{w})$.
@@ IV-B Complexity of OSD4
IV-C Reliable Subset Reduction combined with OSD
As introduced in Section III-B, the RSR procedure removes highly reliable error variables from the parity-check constraints, producing a reduced linear system. This reduced system can then be solved using OSD or any other suitable combinatorial decoder to estimate the remaining unreliable variables.
Increasing the OSD order $w$ generally improves decoder performance. In the extreme case, when $w=n+k$ and $\delta=n$, OSD4-$(n+k)$ with $n$-ADD achieves maximum likelihood decoding. However, larger $w$ comes at the cost of increased computational complexity.
By reducing the problem size through RSR, we can lower the computational overhead in OSD, particularly at low error rates, as demonstrated in Section VII. Moreover, the reduced effective length allows the use of higher-order OSD under a fixed candidate budget. For example, consider a candidate limit of $\Gamma$. For the original problem, OSD4-$2$ tests
@@ IV-C Reliable Subset Reduction combined with OSD
If $v$ highly reliable bits are removed via RSR, the reduced system has 2nv2n-v variables and rank $n-k-(m-m^{\prime})$. The maximum OSD order $w$ that respects the same candidate budget is then
$w=\max\left\{x:\sum_{i=0}^{x}\binom{n+k-v+(m-m^{\prime})}{i}\leq\Gamma\right\}.$ (18)
Even with the same number of candidates, the complexity is lower due to the reduced effective length.
@@ V Approximate Degenerate OSD
In this section, we show how code degeneracy can be exploited to prune the candidate set in OSD. Recall that in OSD4-$w$, a list of $\sum_{i=0}^{w}\binom{n+k}{i}$ candidates is generated by flipping subsets of reliable bits of the OSD4-0 solution and adjusting the unreliable bits accordingly, and the best candidate is selected according to the $\delta$-ADD metric.
In quantum stabilizer codes, however, many of these candidates correspond to degenerate errors, i.e., they differ only by stabilizers and therefore have identical logical effects. Testing such degenerate candidates is unnecessary and can be avoided by tracking their logical syndromes.
Let $\mathbf{L}$} denote a binary logical matrix satisfying ${\mathbf{H}}\Lambda{\mathbf{L}}^{\top}=\mathbf{0},$ so that the logical syndrome of an error $\mathbf{e}$} is given by ${\mathbf{L}}\Lambda{\mathbf{e}}^{\top}.$
Let $\hat{{\mathbf{e}}}$ be the OSD4-0 solution. As in (11), after applying the same column permutation $\pi$, row operations $\rho$, and column permutation $\mu$ used in Gaussian elimination, the logical matrix becomes
${\mathbf{L}}^{\prime}=\mu\!\big(\rho(\pi({\mathbf{L}}\Lambda))\big).$ (19)
Flipping a single reliable bit corresponds to adding the associated column of $\mathbf{L}$$^{\prime}$ to the logical syndrome. Therefore, a reliable bit contributes to a nontrivial logical change if and only if the corresponding column of $\mathbf{L}$$^{\prime}$ is nonzero.
@@
[Degenerate candidate pruning] Consider an $[[n,k,d]]$ stabilizer code whose parity-check matrix has been transformed into the form $\mu(\rho(\pi({\mathbf{H}}\Lambda)))=\begin{bmatrix}{\mathbf{I}}_{n-k}&{\mathbf{A}}\end{bmatrix}$ as in (11), where the last $n+k$ columns correspond to the reliable bits in OSD4-0. Let $\mathbf{L}$$^{\prime}$ be the transformed logical matrix defined above.
@@
    In $n$-ADD OSD, only reliable bits whose corresponding columns in $\mathbf{L}$$^{\prime}$ are nonzero need to be flipped when generating OSD4-$w$ candidates.
@@
    If all columns of $\mathbf{L}$$^{\prime}$ corresponding to the reliable bits are zero, then all OSD4-$w$ candidates are degenerate with the OSD4-0 solution, and OSD4-$w$ reduces to OSD4-0 for any $w$.
@@
Flipping a reliable bit produces a new error candidate whose logical syndrome differs from that of $\hat{{\mathbf{e}}}$ by the corresponding column of $\mathbf{L}$$^{\prime}$. In $n$-ADD, candidates that differ only by stabilizers are equivalent and need not be tested.
1) Therefore, only reliable bits whose columns in $\mathbf{L}$$^{\prime}$ are nonzero can generate candidates in different logical cosets, and only those flips must be considered.
2) If all such columns are zero, then flipping any reliable bit leaves the logical syndrome unchanged, so all OSD4-$w$ candidates are degenerate with $\hat{{\mathbf{e}}}$, and no higher-order OSD can improve upon OSD4-0. ∎
Exploiting Theorem 8-1) with $n$-ADD makes OSD4-$w$ more effective. By incorporating with the RSR algorithm, we can work on a shortened check matrix $\tilde{{\mathbf{H}}}$ and higher order of OSD can be executed. Consequently, (18) can be improved to $w=\operatorname*{arg\,max}_{x}\sum_{i=0}^{x}\binom{u}{i}\leq\Gamma,$ where $u\leq n+k-v+(m-m^{\prime})$ is the number of columns in the shortened check matrix corresponding to nontrivial logical operators.
However, the complexity of $n$-ADD is not practically feasible. We may use $\delta$-ADD by choosing $\delta$ as the highest weight of a set of low-weight stabilizer generators. For the case of a rotated surface code [25], the stabilizer generators are of weight $2$ or $4$, and the next lowest stabilizer weight is $6$. Thus, $n$-ADD can be well approximated by $4$-ADD at error rates lower than $1\%$. Assume that OSD4-0 generates a list of error candidates $\mathcal{E}$. Then, $4$-ADD outputs:
@@
For codes with algebraic structures, the first few lowest-weight terms can be derived, but determining the complete weight enumerator for a general code is an NP-hard problem. As an alternative, we propose selecting a set of $m\geq n-k$ low-weight stabilizer generators and calculating the weight distribution of $m\choose c$} stabilizers, where $c$ is a constant independent of $n$. This method provides a good approximation of the dominant terms for practical purposes. To maintain an overall complexity of $O(n^{3})$, the value of $m\choose c$} should be chosen accordingly.
To achieve low computational complexity, OSD4-0 is often used in practice. Theorem 8 shows that when $n$-ADD is employed, only flips that induce nontrivial logical syndromes need to be considered. However, this argument does not directly apply to OSD4-0. In 0-ADD, candidates are compared solely by their weight. Multiplying an error by a stabilizer may change its weight while leaving its logical coset unchanged, and therefore a degenerate error can become a better candidate under 0-ADD. As a result, in OSD4-0 one cannot restrict bit flips solely based on logical syndromes as in Theorem 8-1). Nevertheless, the degeneracy condition in Theorem 8-2) still applies.
We therefore introduce a weaker but computationally efficient sufficient condition for pruning degenerate candidates.
@@
Let $\tilde{S}\in\{0,1\}^{m^{\prime}\times(2n-v)}$ be a shortened check matrix by the RSR algorithm. Suppose that $\tilde{S}$ can be transformed into $\begin{bmatrix}{\mathbf{I}}&\tilde{{\mathbf{A}}}\end{bmatrix}$ after Gaussian elimination and a column permutation, where $\tilde{{\mathbf{A}}}$ is associated with the reliable bits. If all the columns of $\tilde{{\mathbf{A}}}$ are of weight less than d1d-1, then OSD4-$w$ on the corresponding reduced linear system is equivalent to OSD4-0 for any $w$.
V-A RSR + ADOSD
If many bits are identified as highly reliable, the number of stabilizers associated with these bits would be high. Consequently, the shortened check matrix $\tilde{\mathbf{S}}$ from the RSR algorithm will have fewer rows and lower column weights. Consequently, the sufficient condition in Corollary 9 is satisfied with high probability at low error rates, making higher-order OSD unnecessary.
We combine RSR, degeneracy-aware pruning, and order-adaptive OSD into a unified approximate degenerate OSD (ADOSD) framework, as detailed in Algorithm 4, using 0-ADD. Note that if decoding performance is of higher priority or if the problem size is small enough (e.g., after RSR reduction), $\delta$-ADD can be utilized and can be integrated into Algorithm 4.
@@ V-A RSR + ADOSD
Input: check matrix ${\mathbf{H}}\in\{0,1\}^{m\times 2n}$ of rank nkn-k, syndrome ${\mathbf{s}}\in\{0,1\}^{m\times 1}$, BP output distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, hard-decision reliability vector $\bm{\eta}\in\mathbb{Z}^{n}$, belief distributions $\{{\mathbf{q}}_{i}\}_{i=1}^{n}$, soft reliability threshold $\theta$, hard-decision vector ${\mathbf{e}}\in\{0,1\}^{2n}$, maximum number of OSD flips $\Gamma$, and backup OSD order $w$.
[…]
@@
Apply the RSR algorithm.
if RSR returns failure, then
  apply OSD4-$w$ to the original problem, and return $\hat{{\mathbf{e}}}$;
else RSR returns $\mathbf{s}$$^{\prime}$, $\tilde{{\mathbf{H}}}$, ${\mathbf{e}}^{R}$, and $\sigma$. Suppose $\tilde{{\mathbf{H}}}$ in reduced row echelon form is $I|\tilde{A}$].
  if all the columns of $\tilde{A}$ have weight less than d1d-1, then $w\leftarrow 0$;
@@
[…]
@@
VI Extension to Circuit-Level Noise via Detector Error Models
Our framework is not confined to code-capacity noise; it admits a natural extension to realistic circuit-level noise through detector error models, as we describe below.
At the circuit level, gates, measurements, and ancilla qubits are faulty, and quantum error correction typically involves multiple rounds of stabilizer syndrome extraction circuits. The decoding task is to infer the most likely error configuration given the circuit description and the observed measurement outcomes. This circuit-level decoding problem can be formulated as solving a linear system of equations [60, 70]. Specifically, the error syndrome of a location error is the set of measurement outcomes produced when only that error occurs in the syndrome extraction circuits, while the logical syndrome is given by the commutation relations between the logical operators and the residual data-qubit error after circuit evolution. Collecting the error and logical syndromes of all basis location errors yields a circuit-level parity-check matrix. For an $n$-qubit stabilizer code, the resulting circuit-level decoding problem typically has size $O(n^{3})$ using $O(\sqrt{n})$ syndrome extraction rounds.
STIM [71] provides a detector error model (DEM) for a given syndrome-extraction circuit, which yields a binary circuit-level parity-check matrix ${\mathbf{H}}_{\mathrm{DEM}}$ and a corresponding binary logical matrix ${\mathbf{L}}_{\mathrm{DEM}}$. The error syndrome is the binary vector of detector variables that represent the syndrome difference between two consecutive rounds. Moreover, location errors that produce identical error and logical syndromes are indistinguishable and are therefore merged by STIM into equivalence classes, resulting in a nonuniform probability vector ${\mathbf{p}}_{\mathrm{DEM}}$ that specifies the approximate probability of each error class. Since STIM is a Clifford circuit simulator, it records only commuting observables; in particular, either logical $X$ operators or logical $Z$ operators are monitored. In the following, we assume that logical $X$ operators are observed.
[…]
This abstraction maps a fault-tolerant stabilizer circuit to a binary parity-check matrix with independent error probabilities, which allows our decoder to be applied without modification. For our purposes, this binary decoding problem can be embedded into the quaternary stabilizer formalism by treating ${\mathbf{H}}_{\mathrm{DEM}}$ as $Z$-type stabilizers and ${\mathbf{L}}_{\mathrm{DEM}}$ as logical $Z$ operators, while neglecting commutation relations.
Importantly, circuit-level DEMs are extremely sparse and highly biased: most columns correspond to single-fault events with very small support. As a result, when BP is applied to ${\mathbf{H}}_{\mathrm{DEM}}$, a large fraction of error variables become highly reliable, making RSR particularly effective. In the low-noise regime, RSR typically reduces the effective system size by one to two orders of magnitude, transforming a nominal $O(n^{3})$ circuit-level problem into a reduced system whose size scales roughly linearly with the code distance when error rate is low.
VII Simulation results
We simulate the performance of the MBP4+ADOSD4 scheme on LCS codes [57], BB codes [56], and various 2D topological codes, including toric, rotated surface, color, and twisted XZZX codes [23, 24, 25, 26, 27, 72, 29] (see [82, Table I]) under the code-capacity noise model, and we further simulate rotated surface codes under a circuit-level noise model.